OpenStudy (anonymous):

One runner in a relay race has a 1.50 s lead and is running at a constant speed of 3.25 m/s. The runner has 30.0 m to run before reaching the end of the track. A second runner moves in the same direction as the leader. What constant speed must the second runner maintain in order to catch up to the leader at the end of the race? a. 3.9 m/s c. 2.8 m/s b. 4.75 m/s d. 20 m/s

OpenStudy (anonymous):

@dan815 @Deba_001 @Gravity_Dreams @eric_d @Elsa213 @RLeal @Luigi0210 @Future_UsArmy_MP @Thatkid_marc @iambatman @ikram002p

Elsa213 (elsa213):

yes :)

OpenStudy (anonymous):

=D

OpenStudy (anonymous):

Can you help me out? /: Kind of stuck..

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

@ParthKohli @pinal145 @Abhisar @AnswerMyQuestions @asatt32 @ams98 @davidza6496 @sammixboo @SarahEZZMcK @shkrina @SyedaLovesPie @sheelo_mughal @Stitch_ @zzr0ck3r @ziqbal103 @Zeta @zay813 @CO_oLBoY @CaseyCarns @chrisvillasmil @vksmith7 @victorbell @bloopman @beccamarx19 @bbbco123 @bbblakeee @ninjawolfjj @nuccioreggie @iambatman @OOOPS

OpenStudy (anonymous):



OpenStudy (anonymous):

I am waiting for you to work with.

OpenStudy (anonymous):

okay... so would it be 4.75 m/s?

OpenStudy (anonymous):

no, no, not that simple

OpenStudy (anonymous):

be patient, please. Take steps and confirm if you got me

OpenStudy (anonymous):

Alright.

OpenStudy (anonymous):

the first one race at 3.25m/s so that if he is in front other in 1.5s , then the gap BA = 3.25*1.5 = 4.875 m, ok ?

OpenStudy (anonymous):



OpenStudy (anonymous):

so the total gap the second one have to run is 30+4.875=34.875 m

OpenStudy (anonymous):

One more thing: from A , the time the first one have to spend to get the end is t = \(\dfrac{30}{3.25}=9.23s\) ok?

OpenStudy (anonymous):

Now, with that gap and that time, the second must run at v = \(\dfrac{34.875}{9.23}=3.78m/s\) And he has to catch up the first one before reaching the end, so, the speed should be 3.9m/s

OpenStudy (anonymous):

Oooooh!! Got it! Thank you sooo much!!